# On the randomized Euler scheme for SDEs with integral-form drift

### v1

@@ Abstract.
In this paper, we investigate the problem of strong approximation of the solution of SDEs in the case when the drift coefficient is given in the integral form. Such drift often appears when analyzing stochastic dynamics of optimization procedures in machine learning problems. We discuss connections of the defined randomized Euler approximation scheme with the perturbed version of the stochastic gradient descent (SGD) algorithm. We investigate its upper error bounds, in terms of the discretization parameter $n$ and the size $M$ of the random sample drawn at each step of the algorithm, in different subclasses of coefficients of the underlying SDE. Finally, the results of numerical experiments performed by using GPU architecture are also reported.
@@ 1. Introduction
where $T\in[0,+\infty)$ and $W$ is a $m$-dimensional Wiener process defined on a complete probability space $(\Omega,\Sigma,\mathbb{P})$ with filtration $\{\Sigma_{t}\}_{t\geq 0}$ that satisfies usual conditions. Furthermore, $\eta\in\mathbb{R}^{d}$, $b$ is $\mathbb{R}^{d\times m}$-valued mapping, and $a$ is $\mathbb{R}^{d}$-valued mapping such that
@@
for random element $\xi$ defined on a probability space $(\mathcal{T},\mathcal{A},\mu)$ and the mapping $H:\mathcal{T}\times\mathbb{R}^{d}\to\mathbb{R}^{d}.$ We are interested in the approximation of the value of $X(T).$
@@
SDEs are the natural extension of Ordinary Differential Equations (ODEs). The usage of the Wiener process allows one to incorporate random fluctuations into the trajectories of the solutions, thus making them random and more irregular. Analytical properties, including the existence and uniqueness of solutions of ODEs and SDEs, are respectively investigated in [2] and [6]. For more dedicated examples including delay differential equations (DDEs) and stochastic delay differential equations (SDDEs), the reader is especially referred to [7] and [22].
@@
Such trajectories are able to model data coming from real-life phenomena that include the aforementioned fluctuations. Therefore, SDEs are for example widely used in Finance (see [9, 16] for option pricing, asset pricing, and applications in risk management) and more generally in time series analysis; see Section 11 in [24] for the overview of various estimation techniques for SDE parameters from observed time series data. For more applications, the reader is especially referred to Chapter 7 in [14] and [10, 12, 26].
Regarding the time series analysis, certain classes of ODEs and SDEs is able to model the trajectories of a descenting gradient in the gradient descent algorithms given an optimization problem; see Section 5.2 in [13]. On the other hand, the vast majority of optimization problems that arise from Data Science lead to the equations that contain drift in the aforementioned form (2) because of the random nature of the data. Thus, it makes a natural connection between the point-wise approximation of (1) and the gradient descent algorithms.
@@
In many cases, the existence and uniqueness of the solutions of SDEs are guaranteed but the closed-form formulas are not known. It leads to the usage of numerical schemes for the approximation of trajectories; see [23]. For papers investigating higher-order methods, the reader is referred to Chapter 17 in [11].
@@
In recent years, the properties of so-called randomized Euler algorithms for the approximation of the solutions of differential equations were extensively investigated, which was motivated by the fact that for irregular drift coefficient i.e., discontinuous one w.r.t time variable, classical Euler algorithm fails to converge. See section 3 in [19] for the example that shows a lack of convergence for SDEs in case the drift coefficient is only Borel measurable and the deterministic Euler algorithm is leveraged instead of the randomized one. Currently, miscellaneous randomized numerical schemes are constructed and investigated on various classes of differential equations. For example, in [22] authors investigate the approximation of solutions to SDDEs under Carathéodory-type drift coefficients. On the other hand, research is also focused on point-wise approximation under the assumption that only inexact information about drift and diffusion coefficients is accessible i.e., coefficients are perturbed by some noise; see [17] for the construction of randomized Euler scheme under inexact information and investigation of its error and optimality. In [3], authors show that the randomized two-stage Runge–Kutta scheme is the optimal method among all randomized algorithms on certain classes of ODEs based on standard noisy information. In [4], authors analyze randomized explicit and implicit Euler schemes for ODEs with the right-hand side functions satisfying the Lipschitz condition globally or only locally. On the other hand, in [18] authors deal with point-wise approximation of solutions of scalar SDEs under inexact information, define a randomized derivative-free version of the Milstein algorithm, and investigate its error and optimality. Finally, in [5] authors add randomization to the implicit two-stage Runge-Kutta scheme in order to improve its rate of convergence and analyze its stability in terms of three notions. Similarly for the infinite-dimensional setting, in [21] authors introduce a truncated dimension randomized Euler scheme for SDEs driven by an infinite dimensional Wiener process, with additional jumps generated by a Poisson random measure. Finally, in [20] authors investigate the error of the randomized Milstein algorithm for solving scalar jump-diffusion stochastic differential equations and provide a complete error analysis under substantially weaker assumptions than those known in the literature additionally proving the optimality of the algorithm.
@@
The structure of the paper is as follows. In Section 2 we introduce the considered class of SDEs (1) with admissible coefficients defining the equation. In Section 3, we describe a connection between various gradient descent algorithms and the corresponding variants of Euler schemes. In Section 4, we define a randomized Euler algorithm for the strong approximation of the solutions of (1) and investigate its properties, including its upper $L^{p}$-error bound. Our theoretical results are supported by numerical experiments described in Section 5. Therefore, we also provide the key elements of our current algorithm implementation in Python and CUDA C. In Section 6, we summarize our main findings and list future research. Finally, in Section 6, we state basic facts and provide proofs for the auxiliary lemmas.
@@ 2. Preliminaries
[…]
@@
for $A=[a_{i,j}]_{i,j=1}^{d,m}\in\mathbb{R}^{d\times m}.$ Furthermore, let $0_{d\times m}$ denote a zero-matrix of size $d\times m,$ and let $I_{m}$ denote identity matrix of size $m\times m.$
For $r\in[2,+\infty)$ by the $L^{r}(\Omega)$-norm, either for a random vector or a random matrix, we mean
$\|Y\|_{r}:=\left({\mathbb{E}}\|Y\|^{r}\right)^{1/r}\qquad\mbox{for}\qquad Y:\Omega\to\mathbb{R}^{m}\mbox{ or }Y:\Omega\to\mathbb{R}^{d\times m}.$
Let $T\in[0,+\infty)$ and $p\in[2,+\infty)$. We consider only deterministic initial values, i.e
@@
Let $(\mathcal{T},\mathcal{A},\mu)$ be probabilistic space. A function $a:\mathbb{R}^{d}\to\mathbb{R}^{d}$ belongs to $\mathcal{A}(L,K,p)$ iff:
@@
    the mapping $H:\mathcal{T}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $\mathcal{A}\otimes\mathcal{B}(\mathbb{R}^{d})/\mathcal{B}(\mathbb{R}^{d})$ measurable,
@@
    there exists $K\in[0,+\infty)$ such that $\mathbb{E}\|H(\xi,0)\|^{p}\leq K,$
@@
    there exists mapping $L:\mathcal{T}\to[0,+\infty),$ such that $L\in L^{p}(\mathcal{T},\mathcal{A},\mu)$, and
    $\|H(t,x)-H(t,y)\|\leq L(t)\|x-y\|,$
    for $\mu-$almost all $t\in\mathcal{T}$ and all $x,y\in\mathbb{R}^{d}.$
We further define subclasses of $\mathcal{A}(L,K,p)$ based on function argument dependence. Let $\mathcal{A}_{1}(K,p)$ denote a subclass of $\mathcal{A}(L,K,p)$ such that $a$ depends only on time variable $t$. Similarly, let $\mathcal{A}_{2}(L,p)$ denote a subclass of $\mathcal{A}(L,K,p)$ such that $a$ depends only on spatial variable $x.$
A mapping $b:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ belongs to $\mathcal{B}(K)$ iff:
@@
    there exists $K\in[0,+\infty)$ such that
    $\|b(x)-b(y)\|\leq K\|x-y\|,$
    for all $x,y\in\mathbb{R}^{d}$
We define the general class of admissible tuples $(a,b)$ as follows
$\mathcal{F}(L,K,p)=\mathcal{A}(L,K,p)\times\mathcal{B}(K).$
@@
We use the following notation of asymptotic equalities. For functions $f,g:\mathbb{N}\times\mathbb{N}\to[0,+\infty)$ we write $f(n,M)=\mathcal{O}(g(n,M))$ iff there exist $C>0,n_{0},M_{0}\in\mathbb{N}$ such that for all $n\geq n_{0},\ M\geq M_{0}$ it holds $f(n,M)\leq Cg(n,M)$. Furthermore, we write $f(n,M)=\Omega(g(n,M))$ iff $g(n,M)=\mathcal{O}(f(n,M))$, and $f(n,M)=\Theta(g(n,M))$ iff $f(n,M)=\mathcal{O}(g(n,M))\ \hbox{and}\ f(n,M)=\Omega(g(n,M))$. Unless otherwise stated all constants appearing in estimates and in the "$\mathcal{O}$", "$\Omega$", "$\Theta$" notation will only depend on $\eta,T$ and the parameters of the class $\mathcal{F}(L,K,p)$. Moreover, the same letter might be used to denote different constants.
@@ 3. Connection with gradient descent algorithms
In this section, we describe the connection between various variants of gradient descent algorithms and Euler schemes for the approximation of the solutions of differential equations.
@@
Such systems are called gradient systems. If gradient ${\mathbb{R}^{d}\ni x\mapsto\nabla f(x)\in\mathbb{R}^{d}}$ is continuous and of at most linear growth, then initial-value problem (4) has at least one solution on $[0,T]$ for $T>0$ which is essentially a well-known Peano’s theorem. We refer to Theorem (2.12) on page 252 in [1] where an even more general version of Peano’s theorem was established. If one additionally assumes that $\nabla f$ is locally Lipschitz continuous, then the existence and uniqueness of the solution to (4) can be derived from Theorem 2.2 on pages 104-105 in [8].
Let $x^{*}\in\mathbb{R}^{d}$ be an isolated critical point of $f$ which is also a local minimum. From theorem 4.9 in [2] (see page 142 for the proof), one obtains that if the initial-value $x_{0}$ is sufficiently close to $x^{*},$ then $\lim\limits_{t\to+\infty}x(t)=x^{*}$. Approaching the local minimum $x^{*}$ of $f$ by observing the trajectory $x=x(t)$ while $t\to+\infty$ is sometimes called continuous learning. For $h>0$ and $t_{k}=kh,k=0,1,\ldots,$ the Euler scheme for 4 takes the form
@@
We discussed the origins of the GD method from the perspective of gradient ODEs. We now show how to obtain its stochastic counterpart. Suppose $(\mathcal{T},\mathcal{A},\mu)$ is a probability space and $\xi:\Omega\mapsto\mathcal{T}$ is a random element. We usually assume that real-life data samples are independent realizations of $\xi$ which is further called an apriori distribution of the data. Regarding this random nature of the data, in practice, it usually comes down to the optimization problems which include the expectation formula, i.e
@@
is to be optimized whereas $g$ is usually referred to as a loss function. For example, let $\mathcal{T}=\mathbb{R}^{2},\mathcal{A}=\mathcal{B}(\mathbb{R}^{2}),$ and let $\mu$ denote any distribution on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2})).$ In supervised learning which is one of the branches of machine learning, one’s interest is to find the best mapping $h_{x}(\cdot),$ depending on some parameters $x\in\mathbb{R}^{d},$ such that $h_{x}(\xi_{1})\approx\xi_{2}$ with the usage of algorithms that successively learn from the data. Especially, $h_{x}(\cdot)$ may denote any artificial neural network with weights $x\in\mathbb{R}^{d}.$ In that case, the approximation error can be measured with mean squared error function, i.e loss function of the form $g(\xi_{1},\xi_{2},x)=(h_{x}(\xi_{1})-\xi_{2})^{2}.$ Note that usually only a finite number of data samples can be leveraged. Hence, let $(\xi^{i})_{i=1}^{M}$ denote $M$ independent data samples from the common distribution of $\xi.$ The expectation above is therefore approximated with Monte Carlo sum, i.e
@@
for some $\sigma>0,$ and is called perturbed stochastic gradient descent (PSGD). It turns out that the introduced PSGD scheme (5) is a variant of the randomized Euler scheme for SDEs of the form (1), which is introduced and investigated in the next section of this paper; see [15] where the case of fractional Wiener noise case was considered.

### v2

@@ Abstract.
In this paper, we investigate the problem of strong approximation of the solutions of stochastic differential equations (SDEs) when the drift coefficient is given in integral form. We investigate its upper error bounds, in terms of the discretization parameter $n$ and the size $M$ of the random sample drawn at each step of the algorithm, in different subclasses of coefficients of the underlying SDE presenting various rates of convergence. Integral-form drift often appears when analyzing stochastic dynamics of optimization procedures in machine learning (ML) problems. Hence, we additionally discuss connections of the defined randomized Euler approximation scheme with the perturbed version of the stochastic gradient descent (SGD) algorithm. Finally, the results of numerical experiments performed using GPU architecture are also reported, including a comparison with other popular optimizers used in ML.
@@ 1. Introduction
where $T\in[0,+\infty)$ and $W$ is a $m$-dimensional Wiener process defined on a complete probability space $(\Omega,\Sigma,\mathbb{P})$ with respect to the filtration $\{\Sigma_{t}\}_{t\geq 0}$ that satisfies usual conditions. Furthermore, $\eta\in\mathbb{R}^{d}$, $b$ is $\mathbb{R}^{d\times m}$-valued mapping, and $a$ is $\mathbb{R}^{d}$-valued mapping such that
@@
for the mapping $H:\mathcal{T}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and a random element $\xi$ defined on the probability space $(\Omega,\Sigma,\mathbb{P})$ with values in measurable space $(\mathcal{T},\mathcal{M},\mu)$, where $\mu$ is the law of $\xi$. We assume that the $\sigma$-fields $\sigma(\xi)$ and $\sigma(\bigcup\limits_{t\geq 0}\sigma(W(t)))$ are independent. We are interested in the approximation of the value of $X(T).$
@@
SDEs are the natural extension of Ordinary Differential Equations (ODEs). The usage of the Wiener process allows one to incorporate random fluctuations into the trajectories of the solutions, thus making them random and more irregular. Analytical properties, including the existence and uniqueness of solutions of ODEs and SDEs, are respectively investigated in [barbu] and [cohen_elliott].
@@
Such trajectories are able to model data coming from real-life phenomena that include the aforementioned fluctuations. Therefore, SDEs are for example widely used in Finance (see [glasserman, MERTON] for option pricing, asset pricing, and applications in risk management) and more generally in time series analysis; see Section 11 in [sarkka_solin_2019] for the overview of various estimation techniques for SDE parameters from observed time series data. For more applications, the reader is especially referred to Chapter 7 in [kloeden] and [HARTONO, Iversen, YANG].
Regarding the time series analysis, certain classes of ODEs and SDEs can model the trajectories of a descending gradient in the gradient descent algorithms given an optimization problem; see Section 5.2 in [jentzen2023]. On the other hand, the vast majority of optimization problems that arise from Data Science lead to the equations that contain drift in the aforementioned form (2) because of the random nature of the data; see [pmlr-v40-Ge15]. Furthermore, it turns out that approximating the well-known Stochastic Gradient Descent (SGD) algorithm as SDE allows one to benefit from studying a continuous optimization trajectory while carefully preserving the stochasticity of SGD. The broad overview of the most popular gradient descent algorithms is provided in [ruder2017overviewgradientdescentoptimization] including the aforementioned SGD algorithm. For example, see [Laborde2019NesterovsMW] and [NIPS2014_f0969691] for the direct connection of the Nesterov optimization algorithm with ODEs. For the relevant literature connecting SGD with SDEs, the reader is especially referred to the recent advances provided in [Latz, li2021validitymodelingsgdstochastic, malladi2024sdesscalingrulesadaptive, sirignano2017stochasticgradientdescentcontinuous].
Nonetheless, in [dauphin2014identifyingattackingsaddlepoint], authors claim that the classical SGD algorithm is vulnerable to stucking in saddle points or at least being slowed down by them. In recent years new variants of SGD algorithms were studied to mitigate the negative impact of the occurrence of saddle points, thus resulting in the introduction of the so-called perturbed stochastic gradient descent (PSGD) algorithm. For the relevant literature on the PSGD algorithm, the reader is especially referred to [pmlr-v80-daneshmand18a, du2017gradientdescentexponentialtime, pmlr-v40-Ge15, pmlr-v70-jin17a, pmlr-v75-jin18a, NEURIPS2018_217e342f].
@@
In many cases, the existence and uniqueness of the solutions of SDEs are guaranteed but the closed-form formulas are not known. It leads to the usage of numerical schemes for the approximation of trajectories; see [sauer]. For papers investigating higher-order methods, the reader is referred to Chapter 17 in [kloeden2].
@@
In recent years, the properties of so-called randomized Euler algorithms for the approximation of the solutions of differential equations were extensively investigated, which was motivated by the fact that for irregular drift coefficient i.e., discontinuous one w.r.t time variable, classical Euler algorithm fails to converge. See section 3 in [PRZYB2014] for the example that shows a lack of convergence for SDEs in case the drift coefficient is only Borel measurable and the deterministic Euler algorithm is leveraged instead of the randomized one. Currently, miscellaneous randomized numerical schemes are constructed and investigated on various classes of differential equations. For example, in [PRZYBYLOWICZ2024143] authors investigate the approximation of solutions to SDDEs under Carathéodory-type drift coefficients. On the other hand, research is also focused on point-wise approximation under the assumption that only inexact information about drift and diffusion coefficients is accessible i.e., coefficients are perturbed by some noise; see [morkisz2017] for the construction of randomized Euler scheme under inexact information and investigation of its error and optimality. In [BOCHACIK2021101554], authors show that the randomized two-stage Runge–Kutta scheme is the optimal method among all randomized algorithms on certain classes of ODEs based on standard noisy information. In [BOCHACIK2022], authors analyze randomized explicit and implicit Euler schemes for ODEs with the right-hand side functions satisfying the Lipschitz condition globally or only locally. Finally, in [bochacik2024convergence] authors add randomization to the implicit two-stage Runge-Kutta scheme to improve its rate of convergence and analyze its stability in terms of three notions. Similarly for the infinite-dimensional setting, in [pprzyb2022] authors introduce a truncated dimension randomized Euler scheme for SDEs driven by an infinite dimensional Wiener process, with additional jumps generated by a Poisson random measure. Finally, in [pprzyb2024] authors investigate the error of the randomized Milstein algorithm for solving scalar jump-diffusion stochastic differential equations and provide a complete error analysis under substantially weaker assumptions than those known in the literature additionally proving the optimality of the algorithm.
@@
The structure of the paper is as follows. In Section 2 we introduce the considered class of SDEs (1) with admissible coefficients defining the equation. In Section 3, we describe a connection between various gradient descent algorithms and the corresponding variants of Euler schemes especially describing the connection between randomized Euler scheme for SDEs and PSGD algorithm. We stress that this section does not provide any new theoretical findings in terms of optimization but rather points out interesting relations between these two popular topics. In Section 4, we define a randomized Euler algorithm for the strong approximation of the solutions of (1) and investigate its properties, including its upper $L^{p}$-error bound. Our theoretical results are supported by numerical experiments described in Section 5. Therefore, we also provide the key elements of our current CUDA C algorithm implementation. As a byproduct, in numerical experiments, we also show that the method can be used to search for the local minima. Nevertheless, there are no strict theoretical results on finding the actual minima. In Section 6, we summarize our main findings and list future research. Finally, in Section 6, we state basic facts and provide proofs for the auxiliary lemmas.
@@ 2. Preliminaries
We denote by $\mathbb{N}=\{1,2,\ldots\}$. Let $W=\{W(t)\}_{t\geq 0}$ be a standard $m$-dimensional Wiener process defined on a complete probability space $(\Omega,\Sigma,\mathbb{P})$. By $\{\Sigma_{t}\}_{t\geq 0}$ we denote a filtration, satisfying the usual conditions, such that $W$ is a Wiener process with respect to $\{\Sigma_{t}\}_{t\geq 0}$. We set $\Sigma_{\infty}=\sigma\Bigl(\bigcup_{t\geq 0}\Sigma_{t}\Bigr)$. Next, let $\xi$ be the random element which is defined on the same probability space $(\Omega,\Sigma,\mathbb{P})$ with values in measurable space $(\mathcal{T},\mathcal{M},\mu)$, where $\mu$ is the law of $\xi$. We assume that the $\sigma$-fields $\sigma(\xi)$ and $\sigma(\bigcup\limits_{t\geq 0}\sigma(W(t)))$ are independent.
By $\|\cdot\|$ we denote Frobenius norm for $\mathbb{R}^{d\times m}$ matrices i.e.,
@@
where $A=[a_{i,j}]_{i,j=1}^{d,m}\in\mathbb{R}^{d\times m}.$ Furthermore, let $0_{d\times m}$ denote a zero-matrix of size $d\times m,$ and let $I_{m}$ denote identity matrix of size $m\times m.$ Let $p\in[2,+\infty).$ By $L^{p}(\Omega)$-norm for either a random vector or a random matrix, we mean
$\|Y\|_{p}:=\left({\mathbb{E}}\|Y\|^{p}\right)^{1/p}\qquad\mbox{where}\qquad Y:\Omega\to\mathbb{R}^{m}\mbox{ or }Y:\Omega\to\mathbb{R}^{d\times m}.$
For $T\in[0,+\infty)$, we are interested in strong approximation of the solutions $(X(t))_{t\in[0,T]}$ of (1). The coefficients of the equation are assumed to satisfy the following conditions.
First, we consider only deterministic initial values i.e.,
@@
Next, let $K\in[0,+\infty)$ and let $L:\mathcal{T}\to[0,+\infty)$ be the mapping such that $L\in L^{p}(\mathcal{T},\mathcal{M},\mu)$. Similarly we extend previous notation to $\|L\|_{1}:=\int\limits_{\mathcal{T}}L(t)\mu(\,{\mathrm{d}}t)$ and $\|L\|_{p}:=\|L\|_{{L^{P}(\mathcal{T},\mathcal{M},\mu)}}$. The mapping $H:\mathcal{T}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ belongs to $\mathcal{A}(L,K,p)$ iff:
@@
    $H$ is $\mathcal{M}\otimes\mathcal{B}(\mathbb{R}^{d})/\mathcal{B}(\mathbb{R}^{d})$ measurable,
@@
    $\mathbb{E}\|H(\xi,0)\|^{p}\leq K,$
@@
    $\|H(t,x)-H(t,y)\|\leq L(t)\|x-y\|,$ for $\mu$-almost all $t\in\mathcal{T}$ and all $x,y\in\mathbb{R}^{d}.$
We further define subclasses of $\mathcal{A}(L,K,p)$ based on argument dependence of $H$. Let $\mathcal{A}_{1}(K,p)$ denote a subclass of $\mathcal{A}(L,K,p)$ such that $H$ depends only on time variable $t.$ The strict definition of this subclass is
$\mathcal{A}_{1}(L,K,p):=\Big\{H\in\mathcal{A}(L,K,p):H(t,x)=H(t,0)\ \hbox{for}\ \mu\hbox{-almost all}\ t\in\mathcal{T}\ \hbox{and all}\ x\in\mathbb{R}^{d}\Big\}.$
Similarly, let $\mathcal{A}_{2}(L,p)$ denote a subclass of $\mathcal{A}(L,K,p)$ such that $H$ depends only on the spatial variable $x.$ In a similar manner, this subclass can be defined as
$\mathcal{A}_{2}(L,K,p):=\Big\{H\in\mathcal{A}(L,K,p):H(t,x)=H(0,x)\ \hbox{for}\ \mu\hbox{-almost all}\ t\in\mathcal{T}\ \hbox{and all}\ x\in\mathbb{R}^{d}\Big\}.$
The admissible drift functions $a:\mathbb{R}^{d}\to\mathbb{R}^{d}$ are assumed to be of the form
$a(x)=\mathbb{E}(H(\xi,x))$
where $H$ belongs to $\mathcal{A}(L,K,p)$ or more strictly one of its subclasses. A mapping $b:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ belongs to the class of Lipschitz continuous functions $\operatorname{Lip}(K)$ iff:
@@
    $\|b(x)-b(y)\|\leq K\|x-y\|,$ for all $x,y\in\mathbb{R}^{d}.$
We define the general class of admissible tuples $(H,b)$ as follows
$\mathcal{F}(L,K,p)=\mathcal{A}(L,K,p)\times\operatorname{Lip}(K).$
@@
Finally, we use the following notation of asymptotic equalities. For functions $f,g:\mathbb{N}\times\mathbb{N}\to[0,+\infty)$ we write $f(n,M)=\mathcal{O}(g(n,M))$ iff there exist $C>0,n_{0},M_{0}\in\mathbb{N}$ such that for all $n\geq n_{0},\ M\geq M_{0}$ it holds $f(n,M)\leq Cg(n,M)$. Furthermore, we write $f(n,M)=\Omega(g(n,M))$ iff $g(n,M)=\mathcal{O}(f(n,M))$, and $f(n,M)=\Theta(g(n,M))$ iff $f(n,M)=\mathcal{O}(g(n,M))\ \hbox{and}\ f(n,M)=\Omega(g(n,M))$. Unless otherwise stated all constants appearing in estimates and in the $\mathcal{O}$", $\Omega$", $\Theta$" notation will only depend on $\eta,T$ and the parameters of the class $\mathcal{F}(L,K,p)$. Moreover, the same letter might be used to denote different constants.
@@ 3. Connection with gradient descent algorithms
In this section, we describe the connection between various variants of gradient descent algorithms and Euler schemes for the approximation of the solutions of differential equations. In particular, the stochastic gradient descent algorithms are related to Robbins–Monro method, so readers are especially referred to [Robbins1951ASA]. On the other hand, we would like to stress that the aim of this chapter is not to thoroughly review or prove any theoretical results in the context of gradient descent algorithms, but rather point out that there is a connection with strong approximation of SDEs. For instance, this connection may lead to interesting example equations that originate from Machine Learning.
@@
Such systems are called gradient systems. See [9992691] for more examples including ODE that models Nesterov’s accelerated gradient optimization algorithm. If gradient $\mathbb{R}^{d}\ni x\mapsto\nabla f(x)\in\mathbb{R}^{d}$} is continuous and of at most linear growth, then initial-value problem (4) has at least one solution on $[0,T]$ for $T>0$ which is essentially a well-known Peano’s theorem. We refer to Theorem (2.12) on page 252 in [andres] where an even more general version of Peano’s theorem was established. If one additionally assumes that $\nabla f$ is locally Lipschitz continuous, then the existence and uniqueness of the solution to (4) can be derived from Theorem 2.2 on pages 104-105 in [friedman].
Let $x^{*}\in\mathbb{R}^{d}$ be an isolated critical point of $f$ which is also a local minimum. From Theorem 4.9 in [barbu] (see page 142 for the proof), one obtains that if the initial-value $x_{0}$ is sufficiently close to $x^{*},$ then $\lim\limits_{t\to+\infty}x(t)=x^{*}$. Approaching the local minimum $x^{*}$ of $f$ by observing the trajectory $x=x(t)$ while $t\to+\infty$ is sometimes called continuous learning. For $h>0$ and $t_{k}=kh,k=0,1,\ldots,$ the Euler scheme for 4 takes the form
@@
We discussed the origins of the GD method from the perspective of gradient ODEs. We now show how to obtain its stochastic counterpart. Suppose $(\mathcal{T},\mathcal{M},\mu)$ is a probability space and $\xi:\Omega\mapsto\mathcal{T}$ is a random element. We usually assume that real-life data samples are independent realizations of $\xi$ which is further called an apriori distribution of the data. Regarding this random nature of the data, in practice, it usually comes down to the optimization problems which include the expectation formula i.e.,
@@
is to be optimized whereas $g$ is usually referred to as a loss function; see [pmlr-v40-Ge15]. For example, let $\mathcal{T}=\mathbb{R}^{2},\mathcal{M}=\mathcal{B}(\mathbb{R}^{2}),$ and let $\mu$ denote any distribution on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2})).$ In supervised learning which is one of the branches of machine learning, one’s interest is to find the best mapping $h_{x}(\cdot),$ depending on some parameters $x\in\mathbb{R}^{d},$ such that $h_{x}(\xi_{1})\approx\xi_{2}$ with the usage of algorithms that successively learn from the data. Especially, $h_{x}(\cdot)$ may denote any artificial neural network with weights $x\in\mathbb{R}^{d}.$ In that case, the approximation error can be measured with mean squared error function i.e., loss function of the form $g(\xi_{1},\xi_{2},x)=(h_{x}(\xi_{1})-\xi_{2})^{2}.$ Note that usually only a finite number of data samples can be leveraged. Hence, let $(\xi^{i})_{i=1}^{M}$ denote $M$ independent data samples from the common distribution of $\xi.$ The expectation above is therefore approximated with Monte Carlo sum i.e.,
@@
for some $\sigma>0,$ and is called perturbed stochastic gradient descent (PSGD). It turns out that the introduced PSGD scheme (5) is a variant of the randomized Euler scheme for SDEs of the form (1), which is introduced and investigated in the next section of this paper; see [lucchi2022theoretical] where the case of fractional Wiener noise case was considered.